• In case most of you have not noticed, I made changes to the donation amounts as well as the durations for all of the paid membership levels. Things are winding down here, and this seems to be the most prudent way to deal with the "end of life" period of this site.

    Contributor level members - $10 per 30 days
    Sponsors - Eliminated
    Benefactor level members - $25 per 30 days
    Endowment level members - $50 per 30 days

    Account Upgrades Page
  • Responding to email notices you receive.
    **************************************************
    In short, DON'T! Email notices are to ONLY alert you of a reply to your private message or your ad on this site. Replying to the email just wastes your time as it goes NOWHERE, and probably pisses off the person you thought you replied to when they think you just ignored them. So instead of complaining to me about your messages not being replied to from this site via email, please READ that email notice that plainly states what you need to do in order to reply to who you are trying to converse with.

New ball pythons morphs/mutations

I got SPAMMED with that garbage too. I'm surprised anyone would pay it any attention considering anyone who must resort to SPAM is not worth a second of my time nor anyone elses.
 
Yahoo!

Second clutch from "Silver ghost" laid today with "Chicago girl". Just as a curiosity this girl had probed 6 and 7 scales on each side. She was almost culled from the collection until I decided to try the ultimate sexing, put the animal with a male! It works!

Regards.
 
Here are the eggs! The moist perlite is actually under the mesh. This technique has given me invariably good results.

Regards.
 

Attachments

  • eggs.jpg
    eggs.jpg
    24.7 KB · Views: 120
Just want to go back a few post again lol !

The Spider is Dominant and not co-dom. There is no difference between the het form and the homo form (super).

Markus Jayne in Canada got that little baby
It's a Super Pastel Pewter.
I hope he doesn't mind that I post picture
spp1.jpg

A side with a pewter pastel
pew-sppw.jpg
 
Don't know if any one saw Ralph Davis website, but he is having clutch from Double Het Lavender Albino Piebald.
So he have 1/16 of chance to hit it. He did the samething last year but didn'T get it. Now he have more clutch, hope he get it :).
 
Good thinking breeding your two tone to other morphs first

<snip>The two tone has been bred this year for the first time. I bred her with a couple males. A hypo and a pastel. This way if she proves to not be dominant (which would be an unlikely outcome in my opinion anyhow) I am not raising babies of no potential to breed back. If she produces said:
at least they are pastels or het hypos that I would raise for future breeding anyway[/B]. That assures me that even if the two tone project dead ends in failure in a few years I did not watse the effort of raising normals.
(emphasis added)

We used the same reasoning when we bred our Bullseye (reproducable) to Pastels. The Bullseye trait appears to be dominant but in itself is not all that flashy. It makes the Pastels look great though. A Bullseye to Bullseye cross increases the white, but still is not a knock out snake on it's own. They are useful for making other morphs prettier.

The bullseye white spot is not unheard of. Our original female just happened to pass it on to 100% of her offspring clutch after clutch. Now we have bred it into a Graziani Pastel and it makes the Pastel hold it's color better and increases the white and yellow.
 
So how about a bunch of "I told you so credits" for me now that the pastel pied has hatched? C'mon. I told you guys the pied was just a simple recessive like others and that combo morphs would happen. A friend of mine may have albino pieds this season.
 
Evan

Remember that we were mainly talking about recessive (such as albino pieds) and not co-dominant traits. These might still manifest themselves in the non-white portion of the pied. Who produced it? One way or the other it will be one of those instances where one feels great satisfaction by being proved wrong!
 
The BoidSmith said:
Evan

Remember that we were mainly talking about recessive (such as albino pieds) and not co-dominant traits. These might still manifest themselves in the non-white portion of the pied. Who produced it? One way or the other it will be one of those instances where one feels great satisfaction by being proved wrong!

Just wanted to give a thumbs up to the logic.

As far as the albino pieds, we will be trying to produce some also; but, there is some theory behind why it would not work. The piebaldism is dependant upon the interferance of pre-melanin substances not being trasfered through mast cells to the skin (you genetisists please correct any errors). If the amelanistic trait removes the ability to produce melanin there would be no melanin to transfer to the island sections of the piebald and no pre-melanin substances to be interfered with as far as migration to the skin. So it might not happen as one might think that there would be albino islands on a white snake. Until we see one, we don't know. We will be trying ourselves knowing it may cancel the piebaldism. This is just my .02 playing devils advocate. Those far more informed would have better arguments than I for either side. We are giving it a shot. We think decades down the road though. If it works than you would want to be on the front side. If it doesn't, we have other projects like most people that are going to take a few generations to see what is possible. The offspring would not be worthless if it did not prove out. Perhaps the two are incompatable but the two traits may show up side by side in the same clutch, just not together on the same animal. We didn't get to see such outstanding animals that we have seen over the last 5 years unless someone took the risks. I applaud the risk takers.
 
evansnakes said:
Even if you get 10-12 eggs the odds are statistically applied per egg in genetics, not by percentage, so it is like gambling.

Although I agree with much of what you've said in this thread I have to disagree with that statement. It is very relevant how many eggs are produced, and the odds most certainly improve with each egg. It is true that already having hatched out those eggs unsuccessfully will not improve future chances of hatching the morph combination successfully, but for all unhatched eggs the number of them most certainly affects the odds. Seems like you have the fallacy of conditional probability in reverse.

At 11 eggs, you already have a better than 50% chance at producing an albino pied. I just recently made a few posts on the mathematical probabilities of producing albino x pied as well as basic combinatorics and the binomial distribution in this thread if you care to look into it further:
http://www.reptilescanada.com/forums/showthread.php?t=3014&page=2
 
I'm sorry but you are wrong. You are applying your form of logic to a scientific formula that works on it's own logic. Yes, technically by the odds of just pure chance your odds are better with the more eggs that you get but that is only perception. Genetics are based on the statistical odss of each offspring. Each egg has the same odds of being or not being a mutation animal as the next and is not affected by the outcome of other hatchlings. That is why you can hatch 10 hets out of 10 eggs or 10 mutations out of 10 eggs.

The albino gene is a simple recessive gene and it makes no difference being bred to a pied or a hypo. You are under a misconception that there is more to the pied gene than there is. It is just a simple recessive mutation. The albino pied is producable but in my opinion is a terrible project. The very extreme color separation that makes pieds amazing will be lost in the double homozygous form and it will just be a white and yellow albino. Not exciting. Much like the dissapointment many people had with the hypo-albino ball python.
 
evansnakes said:
I'm sorry but you are wrong. You are applying your form of logic to a scientific formula that works on it's own logic. Yes, technically by the odds of just pure chance your odds are better with the more eggs that you get but that is only perception.
This is not my form of logic; this is how probability and statistics works.

Genetics are based on the statistical odss of each offspring. Each egg has the same odds of being or not being a mutation animal as the next and is not affected by the outcome of other hatchlings. That is why you can hatch 10 hets out of 10 eggs or 10 mutations out of 10 eggs.
Genetics is governed by the same principles of probability and statistics as anything else. You are perfectly correct that each egg has the same odds; this is what's known as independent trials. 10 hets out of 10 eggs or 10 mutations out of 10 eggs are both outcomes accounted for by probability and statistics. Are these outcomes as likely to occur as 5 eggs hatching and 5 eggs hatching out as mutations? Definitely not. There is only one possible combination each for the outcomes you've described while there are 252 ways to hatch out 5 of each. Hence, that would be a much more likely outcome, and would happen

The albino gene is a simple recessive gene and it makes no difference being bred to a pied or a hypo. You are under a misconception that there is more to the pied gene than there is. It is just a simple recessive mutation.
I haven't suggested anything of the sort. In fact, my calculations are based on the assumption that there's isn't any difference as I am using 1/16 odds when breeding double hets together.

It just seems to me that you've confused the fallacies of independent trials and conditional probability. It is true that the outcome of one egg has absolutely no bearing on the outcome of another egg. It is a common fallacy that past failures increase the odds of future success, and that I am not denying. However, if none of the outcomes were yet known, the odds of success given both past trials and future trials would be higher than the future trials alone.

If you're still not convinced, pick up a book on probability and statistics. I was a math minor in undergrad and I am very certain about what I'm asserting here.
 
Han, I am a bit confused why you would attribute my comment about hypo, pied and albinos being the same thing to your comments when it was in response to the comments made by others stating that the pied gene could be combined with a dominant morph like a pastel but not being compatible with the albino gene. It will be. It is the same exact thing just a different breeding required. As to your comments, you aggree with me but don't aggree with me. I really am impressed with your credentials as a math major in college but your comments suggest that because of that you know more than everyone else. I am educated both formally and practically in working with reptiles and would be willing to wager that I have worked with, produced more animals, species and mutations than you and have more practical hands on experience in the area than you do and I am telling you that it is not simple probability and statistics and that the odds of Medel's genetics are based on each individual egg and not on the number of eggs produced. If you say, as I have stated twice previously, that from 10 eggs you would be more likely to produce albinos from a het x het breeding than if you had 5 eggs, that is just common sense and basic logic. BUT, with genetics you are actually just as likely to get 1 albino as you are to get 9 and if you absorb and sorty a huge amount of data you would realize that.

Call Ralph Davis, Brian Barcyzk, Mark Bell, Tracy Barker or any other breeder who has been doing this for a very long time and they will tell you about the many many times that the outcome defied the odds both good and bad in extreme ways. This is because the odds are per each egg individually and not based on clutch size. They go hand in hand but are two different factors. Argue all you want, I have watched it in my incubators for years and years now.
 
What I can't wait for is the Lucy and Pied pairing. That would look really cool. An all white Pied......


...... and for thse wondering, yes, that is a joke.



Evan you are correct. "They go hand in hand but are two different factors." Good way to put it.
 
evansnakes said:
www.ralphdavisreptiles.com click on the journal to see the day by day record or the birthing record page to view his season at a glance


a couple more new ball pythons this week. there has never been an animal like the ball python when it comes to genetic mutations and potential future genetic projects. E
OMFG! I haven't seen so much platinum since the last BET awards show!
 
Evan and Han,

You both bring up excellent points. When crossing two heterozygous for albino ball pythons each egg has ¼ (25%) chance of being albino, ¼ (25%) chance of being normal, and ½ (50%) chance of being heterozygous for albino. My genetics and probability are probably a little rusty but I agree with Evan, each egg is an independent event. What I think is different is that in a series of “independent events” (eggs) the probability of an albino to show up every time in a clutch is the product of the probabilities of each individual event (egg). This means that the odds of 9 albinos showing in the same clutch is 0.25 x 0.25 x 0.25 x 0.25 x 0.25 x 0.25 x 0.25 x 0.25 x 0.25 = 0.06 or 6% (compared to 25% for each individual egg). Although it might seem small it’s huge when compared to winning the lottery ;) Statisticians and mathematicians please correct me if I’m wrong!

Regards.
 
Dan, I agree with everything you wrote except one item. According to my calculator, 0.25 to the nineth power is 0.000004 instead of 0.06. That's still better than the odds of hitting the lottary jackpot, though.

And the odds of getting all normal babies and no albinos from the mating of two heterozygous albinos is 0.75 to the nth power, where n is the number of eggs. In other words, for nine eggs, the odds of getting no albinos is 0.75 to the nineth power, which equals 0.075 or 7.5%. And the odds of getting at least one albino is 92.5%
 
OK, OK, you will have to forgive me....I’ve always had trouble multiplying by more than 4 decimals on top of my head... :D

Regards.
 
Back
Top